### v3

@@ 3 Asteroseismology’s unique probing power
This review is focused on applications of asteroseismology to rotating stars while omitting jargon, so as to make it well accessible. Major focus will be put on future prospects and opportunities (cf. Fig. 1). Hence, we will only introduce concepts and equations that are absolutely necessary to keep the text self-contained. The reader is referred to the monographs by Unno et al. (1989) and Smeyers & Van Hoolst (2010) for full derivations of the oscillation equations. Further, we point to Aerts et al. (2010) and Basu & Chaplin (2017) for extensive descriptions of analysis tools and the principles of stellar modelling in the modern era of space asteroseismology.
Here, it suffices to know that non-radial oscillations lead to periodic deviations from the equilibrium structure of the star. Each mode moves the fluid elements away from the position they would have without any seismic activity. Each mode displaces the fluid elements with a periodicity corresponding to the mode’s frequency and its nodes. Since this concerns a perturbation of a 3-dimensional (3D) body, three numbers are required to describe the geometry of each mode. Two of these are called “wavenumbers” and stand for the number and position of nodal lines at the stellar surface with respect to the symmetry axis of the oscillations. In the case of a spherical star in equilibrium that gets perturbed by the modes, these wavenumbers are usually denoted as that characterise the nodal lines of a spherical harmonic function used to describe the angular part of the displacement vector due to the mode. Assignment of the third number, , is connected with the number of nodes in the radial component of the eigenfunction. It is more involved than the cases of and , because it depends on the dominant restoring force of the oscillations (Aerts et al. 2010). For the simplest case of radial () oscillations, the pressure gradient is the restoring force and hence these are acoustic eigenmodes (or p-modes) of the star. The radial mode with the th lowest eigenfrequency has nodes in its displacement vector. Since the centre of the star is a node, nodes occur between the centre and the surface of the star and one speaks of the th overtone, while the radial mode having only a node in the star’s centre is called the fundamental mode. The classifcation of non-radial modes () in terms of their overtone is considerably more complex because additional restoring forces, such as buoyancy, come into play. We refer to Takata (2012) for a detailed mathematical description of mode classification in terms of the overtone , denoted here as and for p- and g-modes of radial order , respectively. We end this part on the mode numbers by pointing out that the white dwarf community has historically denoted the radial order by instead of .
@@ 3.1 Stellar evolution models in equilibrium
Modern asteroseismic modelling applications require numerical solutions of the relevant oscillation equations described in the next subsection. The modelling hence relies on realistic unperturbed background stellar models. Any fluid element inside a rigidly rotating star is characterised by its position vector with respect to the star’s centre and the time coordinate measured since the stellar birth (the so-called zero-age-main-sequence or ZAMS defining time zero). The fluid elements will have to fulfil the equations of stellar structure at any time during the star’s evolution (see the monograph on models of rotating stars by Maeder 2009, for the derivations and solutions of these equations).
@@
Since we wish to introduce and discuss the dynamical properties of stellar oscillations, we limit ourselves to a description of the equation of motion and we focus on a rigidly rotating star for now. Expressed in a frame of reference co-rotating with the star’s rotation vector , where is the unit vector along the rotation axis, the equation of motion reads:
@@
In this equation, is the density, is the pressure, stands for the velocity vector, and is the gravitational potential fulfilling the equation with the gravitational constant. Aside from the two accelerations due to the pressure and gravitational forces, the two terms represent extra accelerations due to the joint effect of any active forces not spelled out explicitly, one term for forces active inside the star and the other one for forces imposed by external sources. An example of is the Lorentz force caused by an internal magnetic field or accelerations due to radiative forces resulting in a dust-driven or line-driven stellar wind. Additionally, extra accelerations may be caused by external forces such as magnetic fields and/or tides due to one or more companions..
@@
For an unperturbed background model of a single star not subject to any extra forces aside from those due to pressure gradients, gravity, and rotation, the equilibrium state at time is described by . In that case, it follows from Eq. (1) that the star is a spheroid flattened by its centrifugal force (see Espinosa Lara & Rieutord 2013, for an extensive discussion). Most of the current stellar evolution codes simplify the flattening of stars caused by rotation and describe the gaseous spheroids by using only one spatial coordinate, for instance the distance from a fluid element to the centre of the star. Thus any fluid element is characterised by two coordinates, denoted here as with the star’s age.
@@
Almost all asteroseismic applications discussed in this review are based on such 1-dimensional (1D) background models evolving with time, assuming either rigid rotation with constant frequency or shellular rotation (Zahn 1992) for which the rotation frequency only depends on , , and not on latitude or longitude . Shellular rotation results from the assumption of strong horizontal turbulence, forcing a constant rotation rate along isobars. We discuss prospects for asteroseismic applications relying on more realistic yet more complex 2D or 3D background models in the last section.
@@ 3.2 The wave equations describing the dynamics of stellar oscillations in a rotating star
Each oscillation mode causes the fluid elements in the star to get displaced from their equilibrium position according to the Lagrangian vector . In general, the oscillation equations for a rotating star are of the form
where a superscript dot stands for a time derivative and the operators group all acting forces such that their application represents a collection of terms of -th order in for . In order to solve Eq. (2), one can take various approximations and approaches, depending on the importance of the Coriolis, centrifugal, magnetic, and tidal forces. The validity of approximations depends strongly on the mode frequency regimes with respect to the rotation frequency, whether or not we have to take into account deformation due to the centrifugal force, and whether we deal with rigid or non-rigid rotation. We briefly discuss some of the options but refer to the references for details.
Asteroseismology is most often applied in a linear framework, where it is assumed that any perturbation of the background equilibrium model stemming from an oscillation mode is sufficiently small to ignore effects of order higher than one in the displacement in Eq. (2). This means that we can ignore all terms on the right-hand side in Eq. (2). In this case, the equation of motion in Eq. (1) due to the mode is the much simpler version of Eq. (2), namely
@@
with a linear function determined by the equilibrium values and 1st-order perturbations of the density, pressure, and gravitational potential (e.g., Aerts et al. 2010, Chapter 3). This is the simplest version of the wave equation to solve for a family of linear spheroidal non-radial oscillation modes. We highlight some of the latest findings of helio- and asteroseismology in this simplest approximation in Sect. 4.1.
@@
Versions of non-linear theory of non-radial oscillations based on Eq. (2) have also been developed, up to second (Dziembowski 1982; Buchler & Goupil 1984), third (Van Hoolst & Smeyers 1993; Buchler et al. 1995; Mourabit & Weinberg 2023) or fourth (Van Hoolst 1994) order in while ignoring the rotation of the star or by considering it to cause a small perturbation only. Yet, applications of asteroseismic modelling based on such higher-order (in ) theory are so far scarce. Applications occur for oscillation modes in white dwarfs (Zong et al. 2016b), subdwarfs (Zong et al. 2016a), red giants (Weinberg & Arras 2019; Weinberg et al. 2021), and a Sct star (Mourabit & Weinberg 2023), all of which treat rotation as a small perturbation for the computation of the displacement vectors. Modelling of tides in close binaries may also require non-linear non-radial oscillation theory to solve Eq. (2) when linear tides do not provide a sufficiently accurate description, as theorised by Weinberg et al. (2012); Ogilvie (2014).
@@
A different aspect of simplifying Eq. (1) and Eq. (2) concerns the treatment of rotation, notably the importance of the Coriolis and centrifugal forces and how they compare to the Lorentz force. The rotational effect in the equation of motion leads to terms up to . The Coriolis force creates vorticity and it may become the dominant restoring force, instead of the pressure force or gravity. Hence one encounters additional families of modes, which do not exist in non-rotating stars. A well-known example from both geophysics and astrophysics concerns the family of toroidal modes. For a pedagogical derivation and discussion of the various types of mode families in rotating stars, we refer to Townsend (2003b).
@@
The strategy to compute oscillation modes in a rotating star depends entirely on how the mode periods compare to the rotation period. If these two are of comparable order, one cannot treat the rotation as a small effect to compute (see also the next section). If the rotation happens on a much slower timescale than the one of the oscillations, one can treat the rotational effects in a perturbative approach. This is typically a good strategy when the rotation period is at least ten times longer than the mode periods. In such a case, perturbation theory for proper computation of can be developed up to any order in , where is considered as a small expansion parameter to compute the displacement in the form . We note that such methodology only makes sense if is sufficiently small as an expansion parameter, hence the modes should have far shorter periods in the frame of reference co-rotating with the star compared to the rotation period. Within such an approach, one must then also decide whether the deformation of the star, which is , can be ignored or not. If it must be taken into account, the seismic modelling requires a way to incorporate the deformation, either at the level of the equilibrium structure, or at the level of the expression for , or both. Often one considers the simplest deformation only at the level of the equilibrium structure, where the contribution to the potential is approximated by its spherically symmetric component (cf. Eq. (30) in Aerts 2021). Whatever the choice of how to deal with the deformation, one also needs to specify if and up to which order the terms due to the Coriolis and centrifugal forces couple to each other when computing .
@@
A variety of 1st-, 2nd-, and 3rd-order perturbative non-radial pulsation theories for the computation of spheroidal and toroidal families of mode solutions is available in the literature, relying on the assumption that all the terms within the operator as well as the term in Eq. (2) cause only small deviations from the solutions to Eq. (5) represented by . We remark that the computation of from perturbation theory to treat the rotation may involve mode coupling, particularly among spheroidal and toroidal modes, but this is a different matter than adopting non-linear oscillation theory by relying on the right-hand side of Eq. (2). The series in perturbation theory to compute linear modes of rotating stars can be truncated after taking sufficient terms in numerical computations, depending on the envisioned precision required for the modelling application. Perturbative approaches for asteroseismic modelling adhering to the requirement were developed theoretically already long before the modern era of space asteroseismology (e.g., Ledoux 1951; Saio 1981; Gough & Thompson 1990; Dziembowski & Goode 1992, 1996; Soufi et al. 1998; Daszyńska-Daszkiewicz et al. 2002; Karami 2008, each of which containing a lot of technical details).
@@
In summary, for each particular asteroseismic modelling application, proper balancing between the acting forces in Eq. (1) and Eq. (2) is necessary, in order to decide upon the best theoretical formalism and the most suitable numerical approach to calculate the oscillation modes. The choice between a perturbative or non-perturbative approach to calculate the modes of equilibrium models depends entirely on the frequency regime of the detected modes under investigation, notably the value of (see e.g., Ballot et al. 2010, 2013). This is the major reason why we organise applications discussed in the next sections according to the regimes of the frequencies of the detected oscillations.
@@ 3.3 Oscillation mode frequency regimes
All stars rotate, even if many do so slowly, by which we mean that their rotation velocity is only a small fraction (say, up to 10%) of their critical break-up velocity. The seemingly simple concept of break-up velocity, defined as the velocity of fluid elements at the stellar equator large enough to overcome the gravitational attraction, is non-trivial. Indeed, the centrifugal acceleration due to fast rotation can be accompanied by extra outward forces making it easier to overcome the inward force of gravity. An example is the case of a radiation-driven wind (Kudritzki & Puls 2000) or the tidal pull by a companion (Fuller et al. 2019), helping matter to escape more easily from the star compared to the case without such helpful forces.
In this review, we focus on intermediate- and high-mass stars with sufficient detected and identified non-radial oscillation modes in terms of their angular mode dependence to scrutinise stellar equilibrium models. That is to say that the identified modes allow for quantitative measurements of the internal rotation profile and possibly the internal magnetic field. With a few exceptions, asteroseismic modelling based on sufficient modes to achieve this in such types of stars has only begun in the space asteroseismology era. At the time of writing, stars fulfilling these requirements have a mass typically below some 25 M for the core-hydrogen burning phase, and lower for the hydrogen-shell or core-helium burning phases. We will therefore ignore the effect of a dynamical wind and rely on a static outer boundary condition to compute the oscillation modes via Eq. (2). We treat the case of tidal forces due to a close companion in Sect. 6.
@@
where is the mass and the equatorial radius of the star, to compare with the star’s oscillation frequencies in a co-rotating frame. The equatorial radius of a star is most often unknown, unless the star has a well-resolved interferometric image such as the fast rotating stars Altair (Domiciano de Souza et al. 2005; Monnier et al. 2007; Bouchaud et al. 2020), Rasalhague (Monnier et al. 2010), Vega (Monnier et al. 2012), or Archenar (Domiciano de Souza et al. 2014a). Lack of knowledge of is one of the reasons why the critical rotation rate adopted in calculations often corresponds to the Roche critical frequency, defined as
@@
with the star’s polar radius. These two expressions are not the same because the polar radius only equals when the actual critical rotation velocity of the star is reached (which is never the case, as the star would no longer exist). For a more elaborate discussion on the relationship between and , we refer to Rieutord et al. (2016). In citing results below, we will explicitly mention the definition adopted by the authors for the critical rotation frequency whenever available.
@@
As of now, we define a star as a slow rotator if its rotation frequency is less than 10% of its Keplerian critical frequency, a moderate rotator if it has a rotation frequency between 10% and 70% of its Keplerian critical frequency, and a fast rotator if its rotation frequency is above 70% of the Keplerian critical frequency. This may differ substantially from the definition adopted by spectroscopists, who usually only consider to decide if a star is a fast or slow rotator. As an illustration of this, we computed and for the 12 M Cep star HD 192575, which has a measured km s and whose oscillation spectrum is shown in Fig. 6 (discussed further in the text). Based on the high-resolution spectroscopic estimate of one would be tempted to classify this star as a slow rotator because most B-type stars have five to fifteen times higher . However, asteroseismologists are not hampered by the unknown factor as they measure directly from the oscillation frequencies (see Pedersen 2022a, for a sample of B stars). For HD 192575 this leads to an equatorial rotation velocity between 75 and 100 km s and an inclination angle between and (Burssens et al. 2023). Moreover, what matters for asteroseismic modelling is the ratio of twice the rotation frequency and the oscillation mode frequency in a frame of reference co-rotating with the star, the so-called mode spin parameter defined as . From the asteroseismic modelling of this star by Burssens et al. (2023), we find that the star’s rotation frequency at the position of the gradient zone adjacent to the receding convective core has and , while its radiative envelope near the surface rotates at and . The spin parameters of the dipole and quadrupole modes range from about 5% to 20%. Hence we classify this star as a moderate rotator because 2nd-order rotational effects should not be neglected to achieve an optimal interpretation of its observed rotationally split mode frequencies shown in Fig. 6. We come back to the unique asteroseismic study of this star as a key example of low-order p and g mode asteroseismology of a supernova progenitor in Sect. 4.2.
@@
with the local sound speed. Pressure modes of degree are only propagative in the region where . Gravity modes, on the other hand, are only propagative if their frequency is below the so-called Brunt-Väisälä frequency , defined as
@@
Finally, some layers in stars may be subject to a magnetic field, whose origin, evolution, and properties vary from star to star. Such layers experience plasma waves restored by the Lorentz force. These waves are characterised by their so-called Alfvén frequency, , corresponding to a wave speed
@@
Fig. 3 assembles all the introduced frequencies into a single axis representing a classification of the types of waves corresponding to their dominant restoring force. The dynamical properties of the resonant eigenmodes depend on the relationship between their frequencies and the Lamb, Brunt-Väisälä, rotation, and Alfvén frequencies. The types of approximations that can be made to compute the eigenmode frequencies via Eq. (3) for are also indicated on the figure. In particular, given that almost all stars rotate with periods shorter than a decade, we make a distinction of their waves in the sub-inertial and super-inertial regimes, defined as those having spin parameter above or below 1. We refer to the extensive Table A.1 in Aerts et al. (2010) for typical values of mode periods for all the different types of pulsators, along with typical ranges of the dominant mode amplitudes, luminosities, and effective temperatures corresponding to their excitation mechanism and position in the HRD discussed in Fig. 1 of Aerts (2021).
@@
We stress here that the frequencies of gravity modes or internal gravity waves of a moderate or fast rotator detected by an observer in the inertial frame of reference characterised by are appreciably different from the frequencies computed for a reference frame co-rotating at frequency with the star characterised by . This is due to purely geometrical reasons, even without taking into account the Coriolis or centrifugal forces. We can easily understand this based on the simplest case of a slow rotator, for which the geometry of an eigenmode can be well approximated by a spherical harmonic, . In terms of time dependence, the transformation from the co-rotating to an inertial reference frame follows from

### v4

@@ 3 Asteroseismology’s unique probing power
This review is focused on applications of asteroseismology to rotating stars, while omitting jargon so as to make it easily accessible. A major focus is put on future prospects and opportunities (see Fig. 1). Hence, we only introduce concepts and equations that are absolutely necessary to keep the text self-contained. We refer to the monographs by Unno et al. (1989) and Smeyers & Van Hoolst (2010) for full derivations of the oscillation equations. Further, we point to Aerts et al. (2010) and Basu & Chaplin (2017) for extensive descriptions of the analysis tools and the principles of stellar modelling in the modern era of space asteroseismology.
Here it is enough to know that non-radial oscillations lead to periodic deviations from the equilibrium structure of the star. Each mode moves the fluid elements away from the position they would have without any seismic activity. Each mode displaces the fluid elements with a periodicity corresponding to the mode’s frequency and its nodes. Since this perturbs a three-dimensional (3D) body, three numbers are required to describe the geometry of each mode. Two of these are called ‘wavenumbers’ and stand for the number and position of nodal lines at the stellar surface with respect to the symmetry axis of the oscillations. In the case of a spherical star in equilibrium that gets perturbed by the modes, these wavenumbers are usually denoted as , which characterise the nodal lines of a spherical harmonic function used to describe the angular part of the displacement vector due to the mode. Assignment of the third number, , is connected with the number of nodes in the radial component of the eigenfunction. It is more involved than the cases of and because it depends on the dominant restoring force of the oscillations (Aerts et al. 2010). For the simplest case of radial () oscillations, the pressure gradient is the restoring force, and hence these are acoustic eigenmodes (or p modes) of the star. The radial mode with the -th lowest eigenfrequency has nodes in its displacement vector. Since the centre of the star is a node, nodes occur between the centre and the surface of the star and one speaks of the -th overtone, while the radial mode having only a node in the star’s centre is called the fundamental mode. The classifcation of non-radial modes () in terms of their overtone is considerably more complex because additional restoring forces, such as buoyancy, come into play. We refer to Takata (2012) for a detailed mathematical description of mode classification in terms of the overtone , denoted here as pn and gn for p and g modes of radial order , respectively. We end this part on the mode numbers by pointing out that the white dwarf community has historically denoted the radial order as instead of .
@@ 3.1 Stellar evolution models in equilibrium
Modern asteroseismic modelling applications require numerical solutions of the relevant oscillation equations described in the next sub-section. The modelling hence relies on realistic unperturbed background stellar models. Any fluid element inside a rigidly rotating star is characterised by its position vector with respect to the star’s centre and the time coordinate measured since the birth of the star (called the zero age main sequence or ZAMS defining time zero). The fluid elements have to fulfil the equations of stellar structure at any time during the star’s evolution (see the monograph on models of rotating stars by Maeder 2009, for the derivations and solutions of these equations).
@@
Since we wish to introduce and discuss the dynamical properties of stellar oscillations, we limit ourselves to a description of the equation of motion, and we focus on a rigidly rotating star for now. Expressed in a frame of reference co-rotating with the star’s rotation vector , where is the unit vector along the rotation axis, the equation of motion reads
@@
In this equation, is the density, is the pressure, is the velocity vector, and is the gravitational potential fulfilling the equation with the gravitational constant. Aside from the two accelerations due to the pressure and gravitational forces, the two terms represent extra accelerations due to the joint effect of any active forces not spelled out explicitly, one term for forces active inside the star and the other one for forces imposed by external sources. An example of is the Lorentz force caused by an internal magnetic field or accelerations due to radiative forces resulting in a dust-driven or line-driven stellar wind. Additionally, extra accelerations may be caused by external forces such as magnetic fields and/or tides due to one or more companions.
@@
For an unperturbed background model of a single star not subject to any extra forces aside from those due to pressure gradients, gravity, and rotation, the equilibrium state at time is described by . In that case, it follows from Eq. (1) that the star is an oblate spheroid flattened due to its centrifugal force (see Espinosa Lara & Rieutord 2013, for an extensive discussion). Most of the current stellar evolution codes simplify the oblateness of stars caused by rotation and describe the gaseous spheroids by using only one spatial coordinate, for instance the distance from a fluid element to the centre of the star. Thus, any fluid element is characterised by two coordinates, denoted here as with the star’s age.
@@
Almost all the asteroseismic applications discussed in this review are based on such one-dimensional (1D) background models evolving with time, assuming either rigid rotation with constant frequency or shellular rotation (Zahn 1992) for which the rotation frequency only depends on and not on latitude or longitude . Shellular rotation is denoted here as and results from the assumption of strong horizontal turbulence, forcing a constant rotation rate along isobars. We discuss prospects for asteroseismic applications relying on more realistic yet more complex 2D or 3D background models in the last section.
@@ 3.2 The wave equations describing the dynamics of stellar oscillations in a rotating star
Each oscillation mode causes the fluid elements in the star to become displaced from their equilibrium position according to the Lagrangian vector . In general, the oscillation equations for a rotating star are of the form
where a superscript dot stands for a time derivative and the operators group all acting forces such that their application represents a collection of terms of -th order in for . In order to solve Eq. (2), one can take various approximations and approaches, depending on the importance of the Coriolis, centrifugal, magnetic, and tidal forces. The validity of approximations depends strongly on the mode frequency regimes with respect to the rotation frequency, whether or not we have to take into account deformation due to the centrifugal force, and whether we are dealing with rigid or non-rigid rotation. We briefly discuss some of the options, but refer to the references for details.
Asteroseismology is most often applied in a linear framework, where it is assumed that any perturbation of the background equilibrium model stemming from an oscillation mode is sufficiently small to ignore the effects of order higher than one in the displacement in Eq. (2). This means that we can ignore all terms on the right-hand side in Eq. (2). In this case, the equation of motion in Eq. (1) due to the mode is the much simpler version of Eq. (2), namely
@@
with a linear function determined by the equilibrium values and first-order perturbations of the density, pressure, and gravitational potential (e.g. Aerts et al. 2010, Chapter 3). This is the simplest version of the wave equation to solve for a family of linear spheroidal non-radial oscillation modes. We highlight some of the latest findings of helio- and asteroseismology in this simplest approximation in Sect. 4.1.
@@
Versions of non-linear theory of non-radial oscillations based on Eq. (2) have also been developed, up to second (Dziembowski 1982; Buchler & Goupil 1984), third (Van Hoolst & Smeyers 1993; Buchler et al. 1995; Mourabit & Weinberg 2023), or fourth (Van Hoolst 1994) order in , while ignoring the rotation of the star or by considering it to cause only a small perturbation. However, applications of asteroseismic modelling based on higher-order (in ) theory are scarce. They occur for oscillation modes in white dwarfs (Zong et al. 2016b), sub-dwarfs (Zong et al. 2016a), red giants (Weinberg & Arras 2019; Weinberg et al. 2021), and a Sct star (Mourabit & Weinberg 2023), all of which treat rotation as a small perturbation for the computation of the displacement vectors. Modelling of tides in close binaries may also require non-linear non-radial oscillation theory to solve Eq. (2) when linear tides do not provide a sufficiently accurate description, as theorised by Weinberg et al. (2012); Ogilvie (2014).
@@
A different aspect of simplifying Eq. (1) and Eq. (2) deals with the treatment of rotation, notably the importance of the Coriolis and centrifugal forces and how they compare to the Lorentz force. The rotational effect in the equation of motion leads to terms up to . The Coriolis force creates vorticity. It may become the dominant restoring force instead of the pressure force or gravity. Hence, one encounters additional families of modes, which do not exist in non-rotating stars. A well-known example from both geophysics and astrophysics involves the family of toroidal modes. For a pedagogical derivation and discussion of the various types of mode families in rotating stars, we refer to Townsend (2003b).
@@
The strategy used to compute oscillation modes in a rotating star depends entirely on how the mode periods compare to the rotation period. If these two are of comparable order, the rotation cannot be treated as a small effect to compute (see also the next section). If the rotation happens on a much longer timescale than that of the oscillations, the rotational effects can be treated in a perturbative approach. This is typically a good strategy when the rotation period is at least ten times longer than the mode periods. In such a case, perturbation theory for proper computation of can be developed up to any order in , where is a small expansion parameter for computing the displacement in the form . We note that this methodology only makes sense if is sufficiently small as an expansion parameter, and hence the modes should have far shorter periods in the frame of reference co-rotating with the star compared to the rotation period. Within such an approach, one must then also decide whether the deformation of the star, which is , can be ignored or not. If it must be taken into account, the seismic modelling requires a way to incorporate the deformation, either at the level of the equilibrium structure or at the level of the expression for , or both. Often one considers the simplest deformation only at the level of the equilibrium structure, where the contribution to the potential is approximated by its spherically symmetric component (see Eq. (30) in Aerts 2021). Whatever the choice of how to deal with the deformation, one also needs to specify whether and up to what order the terms due to the Coriolis and centrifugal forces couple to each other when computing .
@@
A variety of first-, second-, and third-order perturbative non-radial pulsation theories for the computation of spheroidal and toroidal families of mode solutions is available in the literature, relying on the assumption that all the terms within the operator as well as the term in Eq. (2) cause only small deviations from the solutions to Eq. (5) represented by . We note that the computation of from perturbation theory to treat the rotation may involve mode coupling, particularly among spheroidal and toroidal modes, but this is a different matter than adopting non-linear oscillation theory by relying on the right-hand side of Eq. (2). The series in perturbation theory used to compute linear modes of rotating stars can be truncated after taking sufficient terms in numerical computations, depending on the envisioned precision required for the modelling application. Perturbative approaches for asteroseismic modelling adhering to the requirement were developed theoretically long before the modern era of space asteroseismology (e.g. Ledoux 1951; Saio 1981; Gough & Thompson 1990; Dziembowski & Goode 1992, 1996; Soufi et al. 1998; Daszyńska-Daszkiewicz et al. 2002; Karami 2008, each of which containing a lot of technical details).
@@
In summary, for each particular asteroseismic modelling application, proper balancing between the acting forces in Eq. (1) and Eq. (2) is necessary in order to decide upon the best theoretical formalism and the most suitable numerical approach to calculate the oscillation modes. The choice between a perturbative or non-perturbative approach to calculate the modes of equilibrium models depends entirely on the frequency regime of the detected modes under investigation, notably the value of (see e.g. Ballot et al. 2010, 2013). This is the major reason why we organise the applications discussed in the following sections according to the regimes of the frequencies of the detected oscillations.
@@ 3.3 Oscillation mode frequency regimes
All stars rotate, even if many do so slowly, by which we mean that their rotation velocity is only a small fraction (up to 10%) of their critical break-up velocity. The seemingly simple concept of break-up velocity, defined as the velocity of fluid elements at the stellar equator high enough to overcome the gravitational attraction, is non-trivial. The centrifugal acceleration due to fast rotation can be accompanied by extra outward forces, making it easier to overcome the inward force of gravity. An example is the case of a radiation-driven wind (Kudritzki & Puls 2000) or the tidal pull by a companion (Fuller et al. 2019), helping matter to escape more easily from the star compared to the case without such helpful forces.
In this review we focus on intermediate- and high-mass stars with sufficient detected and identified non-radial oscillation modes in terms of their angular mode dependence to scrutinise stellar equilibrium models. This means that the identified modes allow quantitative measurements of the internal rotation profile and possibly the internal magnetic field. With a few exceptions, asteroseismic modelling based on sufficient modes to achieve this in such types of stars has only begun in the space asteroseismology era. At the time of writing, stars fulfilling these requirements have a mass typically below 25 M⊙ for the core-hydrogen burning phase, and lower for the hydrogen-shell or core-helium burning phases. We therefore ignore the effect of a dynamical wind and rely on a static outer boundary condition to compute the oscillation modes via Eq. (2). We treat the case of tidal forces due to a close companion in Sect. 6.
@@
where is the mass and the equatorial radius of the star, in comparison with the star’s oscillation frequencies in a co-rotating frame. The equatorial radius of a star is most often unknown, unless the star has a well-resolved interferometric image such as the fast-rotating stars Altair (Domiciano de Souza et al. 2005; Monnier et al. 2007; Bouchaud et al. 2020), Rasalhague (Monnier et al. 2010), Vega (Monnier et al. 2012), or Archenar (Domiciano de Souza et al. 2014). Lack of knowledge of is one of the reasons why the critical rotation rate adopted in calculations often corresponds to the Roche critical frequency, defined as
@@
with the star’s polar radius. These two expressions are not the same because the polar radius only equals when the actual critical rotation velocity of the star is reached (which is never the case as the star would no longer exist). For a more elaborate discussion on the relationship between and , we refer to Rieutord et al. (2016). In citing the results below, we explicitly mention the definition adopted by the authors for the critical rotation frequency whenever available.
@@
Here we define a star as a slow rotator if its rotation frequency is less than 10% of its Keplerian critical frequency, a moderate rotator if it has a rotation frequency between 10% and 70% of its Keplerian critical frequency, and a fast rotator if its rotation frequency is above 70% of the Keplerian critical frequency. This may differ substantially from the definitions adopted by spectroscopists, who usually only consider to decide whether a star is a fast or a slow rotator. As an illustration of this, we computed and for the 12 M⊙ Cep star HD 192575, which has a measured km s-1 and whose oscillation spectrum is shown in Fig. 6 (discussed further in the text). Based on the high-resolution spectroscopic estimate of , one would be tempted to classify this star as a slow rotator because most B-type stars have 5 to 15 times higher . However, asteroseismologists are not hampered by the unknown factor as they measure directly from the oscillation frequencies (see Pedersen 2022a, for a sample of B stars). For HD 192575 this leads to an equatorial rotation velocity between 75 and 100 km s-1 and an inclination angle between and (Burssens et al. 2023). Moreover, what matters for asteroseismic modelling is the ratio of twice the rotation frequency and the oscillation mode frequency in a frame of reference co-rotating with the star, called the mode spin parameter and defined as . From the asteroseismic modelling of this star by Burssens et al. (2023), we find that the star’s rotation frequency at the position of the -gradient zone adjacent to the receding convective core has and , while its radiative envelope near the surface rotates at and . The spin parameters of the dipole and quadrupole modes range from about 5% to 20%. Hence, we classify this star as a moderate rotator because second-order rotational effects should not be neglected to achieve an optimal interpretation of its observed rotationally split mode frequencies shown in Fig. 6. We come back to the unique asteroseismic study of this star as a key example of low-order p- and g-mode asteroseismology of a supernova progenitor in Sect. 4.2.
@@
with the local sound speed. Pressure modes of degree are only propagative in the region where . Gravity modes, on the other hand, are only propagative if their frequency is below the Brunt-Väisälä frequency , defined as
@@
Finally, some layers in stars may be subject to a magnetic field, whose origin, evolution, and properties vary from star to star. Such layers experience plasma waves restored by the Lorentz force. These waves are characterised by their Alfvén frequency, , corresponding to a wave speed
@@
Figure 3 assembles all the introduced frequencies into a single axis representing a classification of the types of waves corresponding to their dominant restoring force. The dynamical properties of the resonant eigenmodes depend on the relationship between their frequencies and the Lamb, Brunt-Väisälä, rotation, and Alfvén frequencies. The types of approximations that can be made to compute the eigenmode frequencies via Eq. (3) for are also indicated in the figure. In particular, given that almost all stars rotate with periods shorter than a decade, we make a distinction of their waves in the sub-inertial and super-inertial regimes, defined as those having spin parameter above or below 1. Throughout the life of a star, the frequencies indicated in Fig. 3 change appreciably. We refer to the extensive Table A.1 in Aerts et al. (2010) for typical values of mode periods for all the different types of pulsators, along with typical ranges of the dominant mode amplitudes, luminosities, and effective temperatures corresponding to their excitation mechanism and evolutionary stage as indicated by their position in the HRD discussed in Fig. 1 of Aerts (2021).
@@
We emphasise here that the frequencies of gravity modes or internal gravity waves of a moderate or fast rotator detected by an observer in the inertial frame of reference characterised by are appreciably different from the frequencies computed for a reference frame co-rotating at frequency with the star characterised by . This is due to purely geometrical reasons, even without taking into account the Coriolis or centrifugal forces. We can easily understand this based on the simplest case of a slow rotator, for which the geometry of an eigenmode can be approximated by a spherical harmonic, . In terms of time dependence, the transformation from the co-rotating to an inertial reference frame follows from